OpenStudy (anonymous):

Use trigonometric identities to transform the left side of the equation into the right side?

OpenStudy (anonymous):

\[\sin ^{2}\theta - \cos ^{2}\theta = 2\sin ^{2}\theta -1 \]

OpenStudy (anonymous):

wait for my question you do this?

OpenStudy (anonymous):

What?

OpenStudy (anonymous):

22 km? I don't know.

OpenStudy (anonymous):

Use Sin^2 + Cos^2 = 1

OpenStudy (anonymous):

But when it's sin^2 - cos^2, how does that work?

OpenStudy (anonymous):

@estudier

OpenStudy (anonymous):

If Sin^2 + Cos^2 = 1, what does - Cos^2 equal?

OpenStudy (anonymous):

-cos^2 = cos^2?